OpenStudy (anonymous):

write a quadradic equation using the points -10,8 using the from ax^2+bx+c=0

OpenStudy (saifoo.khan):

\[\huge (x +10)(x - 8) \]

OpenStudy (anonymous):

what about this 2x^2+20+10 x=4?

OpenStudy (saifoo.khan):

\[\huge \color\maroon {x^2 +2x - 80} \]

OpenStudy (saifoo.khan):

what about this 2x^2+20+10 x=4? is that another ques?

OpenStudy (anonymous):

no an idiots way of trying to make an equation... lol

OpenStudy (anonymous):

it wouldnt have been what I wanted I was trying

OpenStudy (saifoo.khan):

\[\huge lol-i-gave-u-the-perfect-answer!.....\]

OpenStudy (anonymous):

I got it thank you

OpenStudy (anonymous):

ok a student opens up a [pagein a book the product is 1806 what is the page numbers?

OpenStudy (anonymous):

42 and 43 nvm

OpenStudy (saifoo.khan):

42 * 43 = 1806

OpenStudy (anonymous):

im going to bed ttyl good night

OpenStudy (saifoo.khan):

Bye Christey, Im going to Bed.

OpenStudy (saifoo.khan):

Lol, same here! Bye!

OpenStudy (anonymous):

lol that was funny

OpenStudy (saifoo.khan):

bye, C u later! :)